OpenStudy (anonymous):

On Tuesday the temperature was -12o F. For the next four days the temperature climbed ten degrees each day. What is the temperature on Saturday after the increase?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

28 degrees

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

28 is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

coool

ganeshie8 (ganeshie8):

good job!

OpenStudy (anonymous):

o what about this

OpenStudy (anonymous):

-6/35 times -10/14

ganeshie8 (ganeshie8):

\[\large -\dfrac{6}{35} \times -\dfrac{10}{14}\]

OpenStudy (anonymous):

no -6/35 and -10/14

ganeshie8 (ganeshie8):

like this ? \[\large \left( -\dfrac{6}{35} \right)\times \left(-\dfrac{10}{14} \right) \]

OpenStudy (anonymous):

no like this hold on

OpenStudy (anonymous):